OpenStudy (anonymous):

Multiple choice - what is(are) the solution(s) of log 5x+log(x-1)=2? a.-4,5 b.4 c.5 d.4,5

OpenStudy (mertsj):

\[\log5x(x-1)=2\] \[10^2=5x^2-5x\] \[x^2-x-20=0\] \[(x-5)(x+4)=0\] \[x=5 or x-4\]

Directrix (directrix):

c.5

OpenStudy (mertsj):

Discard -4 because a negative argument is illegal. So x = 5

Directrix (directrix):

log (5x) + log (x-1) =2 log[(5x)(x-1)] = 2 Assuming this is a base 10 log 10^2 = [(5x)(x-1)] 100 = 5x^2 - 5x 5x^2 - 5x - 100 = 0 x^2 - x - 20 = 0 (x - 5) (x + 4) = 0 x = 5 or x =- 4 x = -4 is an extraneious root because it produces the log of a negative number x = 5 --> Answer